OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12

OpenStudy (sleepyhead314):

divide both sides by -3 then the way to break up absolute values is when given |a| = b it becomes a = b and a = -b

OpenStudy (anonymous):

-3(x+7)=-12 -3x-21=-12 -3x=-12+21 -3x=9 x=9 divide -3 x=-3